Here's my​ dilemma, I can accept a ​$ 1500 bill or play a game ten times. For each roll of the single​ die, I win ​$700 for rolling 1 or​ 2; I win ​$400 for rolling​ 3; and I lose ​$600 for rolling​ 4, 5, or 6. Based on the expected​ value, I should accept the ​$1500 bill.
A.
The statement does not make sense because the expected value after ten rolls is ____dollars, which is greater than the value of the current bill.
B.The statement makes sense because the expected value after ten rolls is ___ ​dollars, which is less than the value of the current bill.

Answer by rothauserc(4718)

For a fair die, the Probability(P) of rolling a number is 1/6
:
P (rolling 1 or 2) = 1/6 + 1/6 = 2/6 = 1/3
:
P (rolling a 3) = 1/6
:
P (rolling 4 or 5 or 6) = 3/6 = 1/2
:
Expected value on 1 roll of die is 700(1/3) +400(1/6) -600(1/2) = 0
:
Expected value is linear so if you roll the die 10 times, expected value is 0 * 10 = 0
:
Answer is B
: